# DistFlow: A Fully Distributed RL Framework for Scalable and Efficient LLM Post-Training

### v2

@@ Abstract.
Reinforcement learning (RL) has become the pivotal post-training technique for large language model. Effectively scaling reinforcement learning is now the key to unlocking advanced reasoning capabilities and ensuring safe, goal-aligned behavior in the most powerful LLMs. Mainstream frameworks usually employ a hybrid-controller architecture where a single-controller dispatches the overall execution logic and manages overall data transfer and the multi-controller executes distributed computation. For large-scale reinforcement learning, minor load imbalances can introduce significant bottlenecks, ultimately constraining the scalability of the system.
To address this limitation, we introduce  DistFlow [^1] The open-source implementation is publicly available at https://github.com/sii-research/siiRL., a novel, fully distributed RL framework designed to break scaling barrier. We adopt a multi-controller paradigm that dispatches data transfer and execution tasks to all workers, which eliminates the centralized node. This allows each worker to operate independently, leading to near-linear scalability up to thousands of GPUs and dramatic efficiency gains. Furthermore, our architecture decouples resource configuration from execution logic, allowing each worker to have a unique execution flow, offering significant flexibility for rapid and cost-effective algorithmic experimentation. Extensive experiments show that  DistFlow achieves excellent linear scalability and up to a 7x end-to-end throughput improvement over state-of-the-art (SOTA) frameworks.
@@ 1. Introduction
The large language model (LLM) and vision language model (VLM) development paradigm starts with pretraining on massive datasets to build a foundation model, followed by supervised fine-tuning (SFT) (sft, ) to teach it instruction following. While this process creates a knowledgeable and task-capable model, it inherently falls short in ensuring reliable alignment with human values and robustly performing complex reasoning. Hence, the paradigm incorporates a third stage: Reinforcement Learning (rlhf, ), the key technique responsible for the advanced capabilities of modern AI like Deepseek-R1 (guo2025deepseek, ), GPT-4o (gpt4o, ), Gemini 2.5 Pro (gemini, ), Claude 4 (claude4, ) and Grok 4 (Grok4, ). Unlike earlier stages that rely on static data, RL is a dynamic and goal-directed optimization process. It quantifies human preferences by training a reward model, which serves as a guiding signal to shape the model’s output. This process reinforces advanced reasoning capabilities and enforces alignment with human values.
@@
Figure 1. Popular RL algorithms, specifically (a) Proximal Policy Optimization and (b) Group Relative Policy Optimization, can be modeled as a DAG. Each node in the graph represents a distinct computational operation, such as Actor Generation or Actor Train, while the edges signify the data dependencies that dictate the execution order.
@@
To address those major limitations, we introduce  DistFlow, a fully distributed RL training framework with high throughput efficiency and flexible execution pipeline. By adopting a multi-controller paradigm,  DistFlow eliminates the central node common in traditional frameworks. It distributes data loading, computation, and collection responsibilities evenly across all workers, removing single-node bottlenecks. This decentralized data and computation flow enables the framework to achieve linear scalability up to a thousand-GPU scale and remarkable runtime efficiency.
@@
  • We design and implement DistFlow, a novel RL framework featuring a fully distributed architecture to eliminate the single-node bottleneck. By employing a multi-controller paradigm and decentralized data management, it achieves superior scalability. Its flexible, DAG-defined pipeline also decouples algorithmic logic from physical implementation, enabling rapid experimentation.
  • We conduct extensive evaluations of DistFlow against SOTA systems. Our results demonstrate near-linear scalability up to a thousand-GPU scale and show up to a 7x end-to-end throughput improvement across various algorithms, model sizes, and data-intensive workloads.
@@ 2.1. Reinforcement Learning for LLMs
Models and Workflow. Recent advancements in artificial intelligence have demonstrated that RL provides a powerful framework for enhancing language models beyond their pretrained capabilities (sft, ; rlhf, ), enabling them to better align with human preferences and solve increasingly complex tasks. In the single-step optimization iteration of models trained via large-scale RL, four core models operate in concert (bai2022training, ; touvron2023llama, ): the Actor Model, which serves as the target model for generating responses; the Reward Model, responsible for providing preference-based rewards for specific tasks (stiennon2020learning, ); the Reference Model, which functions as a benchmark for measuring knowledge accuracy; and the Critic Model, which evaluates value. The training process follows a meticulously designed three-step iterative workflow: Generation, Evaluation, and Training. During the Generation phase, the Actor model receives a batch of prompts as input and auto-regressively generates corresponding text responses for each one. Once response generation is complete, the process advances to the Evaluation phase, where the Reward Model, Reference Model, and Critic Model each provide scores for the current responses. The final Training phase constitutes the core of the RL training process, wherein these three scores are integrated to calculate the advantage function, and through backpropagation, only the parameters of the Actor and Critic are updated, while the Reward and Reference remain frozen throughout the process (wang2025reinforcement, ).
Algorithms. Policy Optimization serves as the critical mechanism in reinforcement learning that updates a model’s decision-making strategy to maximize expected rewards by systematically improving the policy based on feedback signals, as shown in Figure 1. PPO enhances stability during training by implementing a clipped objective function that prevents destructive policy updates while maintaining sample efficiency (gao2023scaling, ). GRPO reduces computational costs by eliminating the need for a separate critic model and instead estimates baselines from group scores, making large-scale language model reinforcement learning more efficient (guo2025deepseek, ).
Paradigm Shift. Rather than just predicting the next token, RL transforms language models into agents that can optimize toward specific objectives defined by reward functions (leike2018scalable, ; glaese2022improving, ; zaheer2020big, ). Through this approach, RL frameworks enable models to improve through interaction and feedback, rather than being limited by their initial training data (ziegler2019fine, ; nakano2021webgpt, ). Previously, it was widely believed that pre-training scaling capabilities established the ceiling for model training and inference, but recent research has challenged this view, especially the performance breakthrough of Grok 4 (Grok4, ; raffel2020exploring, ; clark2020electra, ; radford2019language, ). Indeed, while pretraining improvements are showing diminishing returns, RL offers a new avenue for significant advancements in model capabilities (lazaridou2020multi, ; lewis2020retrieval, ; haarnoja2018composable, ).
2.2. Scalability Challenges in RL
Scaling Laws in RL. Scaling brings about emergent model capabilities in advanced AI systems. Additionally, we can observe that after scaling up computing resources, Grok 4 demonstrated emergent abilities in Humanity’s Last Exam tasks, pointing to a promising path for future research. The scaling of computational resources significantly impacts a model’s ability to handle complex tasks, while simultaneously placing higher demands on efficient parallelization strategies to fully leverage these resources (espeholt2019seed, ).
Parallelism Strategies. Common parallel strategies include Data Parallel (DP), Tensor Parallel (TP), Pipeline Parallel (PP), etc. DP generally uses parameter sharding methods like Zero Redundancy Optimizer (ZeRO) (rasley2020deepspeed, ) to parallelize model training across multiple devices, and will provide sample-level parallelism by processing different data batches on different devices while synchronizing gradients. TP, as implemented in Megatron-LM (shoeybi2019megatron, ), decomposes matrix multiplications and distributes them across multiple GPUs, providing operation-level parallelism that effectively addresses memory constraints in large transformer models. PP simply divides the neural network into sequential stages across devices, enabling efficient pipelining of micro-batches through different model layers while reducing peak memory requirements.
Bottleneck of RL scaling While scaling computational resources and data generally improves model performance, largescale RL training faces unique bottlenecks that prevent linear scaling. Unlike supervised learning where more data almost always helps, RL systems can actually experience performance degradation with increased batch sizes, as the correlation between consecutive updates becomes weaker, making learning less stable (espeholt2019seed, ). Simply adding more computational resources doesn’t guarantee proportional performance improvements in RL systems. As RL training progresses, models tend to increasingly overoptimize against reward models, leading to diminishing returns (gao2023scaling, ). This overoptimization creates a fundamental ceiling on performance gains regardless of additional computing power.For instance, the data skewness problem in generation stages causes significant GPU underutilization as only a few long-tailed samples keep expensive computational resources waiting (zhong2024optimizing, ).
@@ 2.3. Distributed System Architectures for RL
Controller Paradigms. Distributed machine learning systems, particularly those designed for reinforcement learning, employ different controller paradigms to manage computation across hardware resources. These paradigms fundamentally shape how algorithms are implemented, scaled, and optimized in practice (espeholt2019seed, ).
Single-Controller. The Single-Controller paradigm employs a centralized controller to manage the overall execution flow of the distributed program. With centralized control logic, users can build core functionalities of the dataflow as a single process, while the controller automatically generates distributed workers to carry out the computation. This approach provides a global view of hardware and dataflow graphs, allowing flexible and optimized resource mapping and execution order coordination among dataflow tasks. However, coordination messages are passed from the controller to all workers, incurring significant dispatch overhead when executing expansive dataflow graphs on large clusters (espeholt2019seed, ).
Multi-Controller. In contrast, the Multi-Controller paradigm distributes control logic by giving each device (or worker) its own controller. This approach is commonly used in RL frameworks (espeholt2019seed, ), similar to recent RL training systems (wang2025reinforcement, ), where multiple long-running distributed programs operate with each component coordinating execution order through hard-coded data synchronization. While this reduces the central coordination bottleneck, it often results in complex implementation and maintenance challenges, particularly when scaling to large model architectures or cluster sizes.
Figure 2. The single-controller bottleneck. All data operations, from initial loading to the circulation of intermediate data between stages, are flowed through the centralized controller, leading to severe communication overhead and scalability limitations.
@@ 2.4. Limitations of existing RL systems
Demanding. To address the efficiency issues of earlier designs, researchers have adopted architectures that rely on a single-controller for high-level orchestration, with some frameworks like verl introducing a hybrid paradigm that merges this with the efficiency of multi-controllers for computation. While this approach can improve system throughput in limited-scale scenarios, its reliance on a single-controller for managing the entire dataflow introduces a fundamental and critical limitation: an inability to scale.
In the current era of AI research, where training frontier models on thousands of GPUs is becoming a core requirement, the ability to scale efficiently is quite important. Systems built on a single-controller dataflow, whether fully centralized or hybrid, are inherently incapable of meeting this demand. This architectural choice, where a centralized node orchestrates all data-related tasks, becomes a severe bottleneck as computational resources increase. This design is overwhelmed by the massive volume of data at large scales, leading to instability and failures that fundamentally constrain research and development. This critical scalability issue, along with an inflexible pipeline design, reveals that such architectures are unsuited for the demands of modern, large-scale AI research.
Single-Controller Bottleneck. As shown in Figure 2, the bottleneck of the single-controller node is particularly serious. This design adopts a single-controller paradigm, where a centralized controller orchestrates the entire dataflow. Its responsibilities encompass all data-related tasks, from the initial loading of the dataset to the large-scale intermediate data across different computational stages. This centralized design forces all intermediate data to flow through controller node, introducing costly ”one-to-all” and ”all-to-one” communication operations. These operations introduce substantial I/O and communication costs, severely impacting system efficiency, especially for data-intensive tasks such as multi-modal or long-text generation. This architecture not only demonstrates inefficiency but fundamentally constrains system scalability. In large-scale distributed training, the peak volume of data can easily overwhelm the controller node’s memory, leading to out-of-memory (OOM) errors and thereby limiting the system’s peak data processing capacity. Consequently, the single-controller acts as a dual bottleneck, constraining both the efficiency and the scalability of the entire system.
@@ 3. DistFlow Overview
Therefore, we introduce DistFlow, a fully distributed RL framework designed for scalability on large-scale clusters. As illustrated in Figure 3, DistFlow employs a multi-controller paradigm that uniformly dispatches all computational and data flow across each GPU.  DistFlow consists of three main components: a DAG Planner (§4), DAG Workers (§5), and a Data Coordinator (§6). The DAG Planner receives a custom DAG file from user and splits the overall DAG into a number of individual DAG tasks, equal to the number of DAG Workers, with each task containing no parallel nodes. The DAG Worker is the core logic executor, with each worker bound to a single GPU; it receives and executes a single DAG Task from the DAG Planner. Each DAG Worker has a Distributed Dataloader to load the corresponding data before task execution. The Data Coordinator manages all data coordination among the DAG Workers. At the end of each stage, it collects data into a series of Distributed Databuffers, which is then dispatched at the start of the following stage.
The workflow of  DistFlow is as follows. Before training, users are required to provide three essential configurations: Model Config, Training Config, and Algorithm Config. When utilizing custom execution pipelines, an additional DAG Config must be specified; however, this configuration becomes optional when employing built-in algorithms such as GRPO or PPO. The Model Config specifies the architecture and the corresponding parallelism strategy for each model in the dataflow. The Training Config and Algorithm Config supply the necessary training settings and hyperparameters, respectively.
Upon receiving these config, the DAG Planner processes the global DAG and decomposes it into the smallest executable units, referred to as DAG Tasks. DAG tasks can be replicated across different DAG Workers. These DAG tasks are then dispatched to a set of DAG Workers, where each DAG Worker is bound to a single GPU, forming the fundamental computational unit. Each DAG Worker initializes by loading its assigned model and data, then setting up the required training and inference engines with the configured parallelization strategy. After initialization, all DAG Workers begin to execute their assigned DAG tasks in parallel. During execution, data need to be redistributed among DAG Workers because the changes in parallelism strategies between different stages. This is managed by the Data Coordinator, which automatically senses the parallelism configuration of each stage. It orchestrates the collection and dispatch of data through a series of Distributed Databuffers, ensuring seamless data flow across the system. The system allocates one databuffer per node.
We implemented our system based on PyTorch (pytorch, ). For resource management of GPU and CPU resources, we use Ray (ray, ) which is an open source framework to build and scale ML and Python applications easily. Our system’s architecture integrates specialized engines for different operations. We use PyTorch Fully Sharded Data Parallel (FSDP) (pytorch, ) as the training engine to execute both inference and training operations. For generation operations, we utilize inference engines such as vLLM (vllm, ) and SGLang (sglang, ), which are designed for efficient auto-regressive generation. To manage these components, and inspired by the hierarchical API design of verl , our system uses the 3DParallelWorker base class.
@@ 4. DAG Planner
Figure 4. Decomposing a user-defined DAG into a sequential execution pipeline. DAG Planner transforms parallel nodes like Inference I and Inference II into a sequential chain.
@@ 4.2. DAG Decomposition
A primary challenge in executing a user-defined DAG is ensuring its safe and efficient adaptation to a resource-constrained, colocated architecture where multiple large models share the same resource pool.
@@
This design decouples the high-level algorithmic logic from the low-level execution plan, allowing researchers to define complex workflows without manually managing serialization. This automated process guarantees that only one node is active at any given time.
@@ 5. DAG Worker
Figure 5. Mapping of node definitions to their corresponding execution functions. The DAG Worker dynamically binds a specific computational function to the node based on node_role and node_type.
@@
This abstracted design is realized through a concrete operational flow. In the Initialization phase, the worker instantiates its environment based on the abstract DAG. It uses a Distributed Dataloader to acquire data, loads the specified models, and initializes backend engines like vLLM, SGLang, PyTorch FSDP, or Megatron. It then materializes the task chain into a concrete execution queue, binding the appropriate function to each node, as illustrated in Figure 5.
During the subsequent Execution phase, the worker enters a loop for each RL iteration. It requests a data batch and sequentially executes each node in the chain. A databuffer serves as an intermediary state manager, providing necessary inputs to each function and storing its outputs. Upon completing the chain, metrics are aggregated at the global rank 0 via an efficient all-to-one communication. The primary benefit of this model is its inherent extensibility. The decoupling of the workflow’s structure (defined in the DAG) from its operational logic (resolved by the function mapping) allows researchers to rapidly innovate. For instance, introducing a new reward calculation method or a different policy loss function does not require altering the core dataflow. A researcher can simply implement the new logic in a custom function and map it to a node in the DAG, seamlessly integrating it into the execution pipeline without modifying the surrounding framework.

### v3

@@ Abstract.
Reinforcement learning (RL) has become the pivotal post-training technique for large language model (LLM). Effectively scaling reinforcement learning is now the key to unlocking advanced reasoning capabilities and ensuring safe, goal-aligned behavior in the most powerful LLMs. Mainstream frameworks usually employ a hybrid-controller architecture where a single-controller dispatches the overall execution logic and manages overall data transfer and the multi-controller executes distributed computation. For large-scale reinforcement learning, minor load imbalances can introduce significant bottlenecks, ultimately constraining the scalability of the system.
To address this limitation, we introduce  DistFlow, a novel, fully distributed RL framework designed to break scaling barrier. We adopt a multi-controller paradigm that dispatches data transfer and execution tasks to all workers, which eliminates the centralized node. This allows each worker to operate independently, leading to near-linear scalability up to  1024 GPUs and dramatic efficiency gains. Furthermore, our architecture decouples resource configuration from execution logic, allowing each worker to have a unique execution flow, offering significant flexibility for rapid and cost-effective algorithmic experimentation. Extensive experiments show that  DistFlow achieves excellent linear scalability and up to a 7x end-to-end throughput improvement in specific scenarios over state-of-the-art (SOTA) frameworks.
@@ 1. Introduction
The large language model (LLM) and vision language model (VLM) development paradigm starts with pretraining (transformer, ) on massive datasets to build a foundation model, followed by supervised fine-tuning (SFT) (sft, ; chung2024scaling, ) to teach it instruction following. While this process creates a knowledgeable and task-capable model, it inherently falls short in ensuring reliable alignment with human values and robustly performing complex reasoning. Hence, the paradigm incorporates a third stage: Reinforcement Learning (rlhf, ), the key technique responsible for the advanced capabilities of modern AI like Deepseek-R1 (deepseek-r1, ), GPT-4o (gpt4o, ), Gemini 2.5 Pro (gemini, ), Claude 4 (claude4, ) and Grok 4 (Grok4, ). Unlike earlier stages that rely on static data, RL is a dynamic and goal-directed optimization process. It quantifies human preferences by training a reward model, which serves as a guiding signal to shape the model’s output. This process reinforces advanced reasoning capabilities and enforces alignment with human values.
@@
Figure 1. Popular RL algorithms, specifically (a) Proximal Policy Optimization and (b) Group Relative Policy Optimization, can be modeled as a DAG.
@@
To address those major limitations, we introduce  DistFlow, a fully distributed RL training framework with high throughput efficiency and flexible execution pipeline. By adopting a multi-controller paradigm,  DistFlow eliminates the central node common in mainstream frameworks. It distributes data loading, computation, and collection responsibilities evenly across all workers, removing single-node bottlenecks. This decentralized data and computation flow enables the framework to achieve linear scalability up to a  1024 GPU scale and remarkable runtime efficiency.
@@
  • We introduce DistFlow, a novel RL framework that achieves high scalability and efficiency through its fully distributed architecture and offers significant flexibility via its DAG-defined design.
  • We conduct extensive evaluations of DistFlow against SOTA systems. Our results demonstrate near-linear scalability up to a  1024 GPU scale and show significant end-to-end throughput improvements across various algorithms, model sizes, and model types, reaching up to 7x in specific scenarios.
@@ 2.1. Reinforcement Learning for LLMs
Models and Workflow. Recent advancements in artificial intelligence have demonstrated that RL provides a powerful framework for enhancing language models beyond their pretrained capabilities (sft, ; rlhf, ), enabling them to better align with human preferences and solve increasingly complex tasks. During a single optimization iteration in large-scale RL, the Actor Model’s update is computed through the interaction of four core models (bai2022training, ; llama2, ): the Actor Model itself generates a response, a Reward Model provides preference-based rewards for this response (stiennon2020learning, ), a Critic Model estimates the expected return (deeprl, ), and a Reference Model imposes a KL-divergence penalty for regularization.
The training process follows a meticulously designed three-step iterative workflow: Generation, Evaluation, and Training. During the Generation phase, the Actor Model receives a batch of prompts as input and auto-regressively generates corresponding text responses for each one. Once response generation is complete, the process advances to the Evaluation phase, where the Reward Model, Reference Model, and Critic Model each provide scores for the current responses. The final Training phase constitutes the core of the RL training process, wherein these three scores are integrated to calculate the advantage function, and through backpropagation, only the parameters of the Actor and Critic are updated, while the Reward and Reference remain frozen throughout the process (roll, ).
Algorithms. Policy Optimization is a fundamental class of RL methods that iteratively refines a model’s policy to maximize the expected cumulative reward based on feedback signals, as illustrated in Figure 1. As a widely-adopted policy optimization algorithm, PPO enhances training stability by employing a clipped surrogate objective function. This mechanism constrains the magnitude of policy updates, thereby preventing destructive changes while maintaining high sample efficiency. Because PPO is reliable and performs well, it has become the accepted standard for fine-tuning large models. GRPO is a variant of PPO that enhances a model’s mathematical reasoning abilities while also optimizing PPO’s memory efficiency. It removes the separate critic model, which uses a lot of computing power, and instead estimates baselines directly from group rewards. This approach trades some baseline accuracy for a large increase in training speed, making it very suitable for large language models that require significant computational resources (deepseek-r1, ; grpo, ).
Paradigm Shift. While pre-training instills broad knowledge, RL introduces a critical paradigm shift from static next-token prediction to dynamic, goal-oriented optimization defined by reward functions (leike2018scalable, ; glaese2022improving, ; zaheer2020big, ). This transition transforms language models into agents that are refined through interaction and feedback, enabling them to improve beyond the limitations of their initial training data (ziegler2019fine, ; nakano2021webgpt, ). As performance gains from scaling pre-training begin to diminish, large-scale RL has emerged as a new frontier for advancing model capabilities (lazaridou2020multi, ; lewis2020retrieval, ; haarnoja2018composable, ). Consequently, applying RL to the largest models has become a core component of the modern development lifecycle, creating an urgent demand for efficient and scalable frameworks to manage its immense computational complexity.
@@ 2.2. Distributed System Architectures for RL
While common parallelism strategies like Data Parallel (DP), Tensor Parallel (TP), and Pipeline Parallel (PP) (deepspeed, ; megatron, ) distribute the workload, the underlying distributed system architecture is critical for orchestrating the complex flow of data and computation. At the heart of this architectural design is the choice of a controller paradigm, which dictates how tasks are managed across hardware resources and fundamentally shapes how algorithms are implemented, scaled, and optimized.
Controller Paradigms. Distributed machine learning systems, particularly those designed for reinforcement learning, employ different controller paradigms (pathways, ) to manage computation across hardware resources. These paradigms fundamentally shape how algorithms are implemented, scaled, and optimized in practice.
Single-Controller. The Single-Controller paradigm employs a centralized controller to manage the overall execution flow of the distributed program. With centralized control logic, users can build core functionalities of the dataflow as a single process, while the controller automatically generates distributed workers to carry out the computation. This approach provides a global view of hardware and dataflow graphs, allowing flexible and optimized resource mapping and execution order coordination among dataflow tasks. However, coordination messages are passed from the controller to all workers, incurring significant dispatch overhead when executing expansive dataflow graphs on large clusters.
Multi-Controller. In contrast, the Multi-Controller paradigm (distml, ) distributes control logic by giving each device (or worker) its own controller. This approach is commonly used in RL frameworks, similar to recent RL training systems (roll, ), where multiple long-running distributed programs operate with each component coordinating execution order through hard-coded data synchronization. While this reduces the central coordination bottleneck, it often results in complex implementation and maintenance challenges, particularly when scaling to large cluster sizes.
Figure 2. The bottleneck of centralized data management on a single controller. All data operations are flowed through the centralized controller, leading to severe communication overhead and scalability limitations.
@@ 2.3. Limitations of existing RL systems
Demanding. In the current era of AI research, training frontier models on thousands of GPUs has become a core requirement. However, systems built on a single-controller dataflow, whether fully centralized or hybrid, are inherently incapable of meeting this demand. This architectural choice creates a severe bottleneck at large scales, leading to instability and failures that fundamentally constrain modern research and development, rendering such designs unsuitable for large-scale AI.
Bottleneck in Single-Controller Paradigm. A critical architectural bottleneck emerges in single-controller paradigm when the central controller node is also tasked with managing the data plane for the entire workflow, as illustrated in Figure 2. In such a design, the controller orchestrates not only the execution flow but also the transfer of all large-scale intermediate data between distributed computational stages. This centralization of the data path forces costly ”one-to-all” and ”all-to-one” communication patterns, introducing substantial I/O and network overhead that severely degrades system efficiency, particularly for data-intensive tasks like multi-modal or long-text generation. Furthermore, this architecture fundamentally constrains system scalability. During large-scale distributed training, the peak volume of intermediate data can overwhelm the controller node’s memory capacity, leading to out-of-memory (OOM) errors and imposing a hard limit on the system’s data processing throughput. Consequently, tasking the single controller with data plane management creates a dual bottleneck, simultaneously limiting both the performance and the maximum scale of the entire system.
@@ 3. DistFlow Overview
Based on these design considerations, we propose DistFlow, a fully distributed RL framework designed for scalability on large-scale clusters. As illustrated in Figure 3, DistFlow employs a multi-controller paradigm that uniformly dispatches all computational and data flow across each GPU. DistFlow consists of three main components: a DAG Planner (§4), DAG Workers (§5), and a Data Coordinator (§6).
At its core, this architecture separates the control flow from the data flow. The DAG Planner translates the user’s high-level DAG into concrete, executable tasks. These tasks are then executed by the DAG Workers, the primary computational units bound to individual GPUs. Concurrently, the Data Coordinator manages the entire data lifecycle, orchestrating the complex data redistribution required when parallelism strategies change between stages. This separation of concerns is critical as it simplifies the overall system logic, allows for independent optimization of computation and data transfer, and provides greater flexibility for complex workflows.
We implemented our system based on PyTorch (pytorch, ). For resource management of GPU and CPU resources, we use Ray (ray, ) which is an open source framework to build and scale ML and Python applications easily. Our system’s architecture integrates specialized engines for different stages. We use PyTorch Fully Sharded Data Parallel (FSDP) (pytorch, ) and Megatron  (megatron, ) as the training engine. For generation stage, we utilize the vLLM (vllm, ) and SGLang (sglang, ) inference engines, which are designed for efficient auto-regressive generation. To manage these components, and inspired by the hierarchical API design of verl , our system uses the 3DParallelWorker base class.
@@ 4. DAG Planner
Figure 4. Decomposing a user-defined DAG into a sequential execution pipeline.
@@ 4.2. DAG Decomposition
A primary challenge in executing a user-defined DAG is ensuring its safe and efficient adaptation to a colocated architecture with limited resources, where multiple large models share the same resource pool.
@@ 5. DAG Worker
Figure 5. Mapping of node definitions to their corresponding execution functions. The DAG Worker dynamically binds a specific computational function to the node based on its attributes.
@@
This abstracted design is realized through a concrete operational flow. In the Initialization phase, the worker instantiates its components based on the abstract DAG. It uses a Distributed Dataloader to acquire data, loads the specified models, and initializes backend engines like vLLM, SGLang, PyTorch FSDP, or Megatron. It then materializes the task chain into a concrete execution queue, binding the appropriate function to each node, as illustrated in Figure 5.
During the subsequent Execution phase, the worker enters a loop for each RL iteration. It requests a data batch and sequentially executes each node in the chain. A databuffer serves as an intermediary state manager, providing necessary inputs to each function and storing its outputs. Upon completing the chain, metrics are aggregated at the global rank 0. The primary benefit of this model is its inherent extensibility. The decoupling of the workflow’s structure (defined in the DAG) from its operational logic (resolved by the function mapping) allows researchers to rapidly innovate. For instance, introducing a new reward calculation method or a different policy loss function does not require altering the core dataflow. A researcher can simply implement the new logic in a custom function and map it to a node in the DAG, seamlessly integrating it into the execution pipeline without modifying the surrounding framework.